OpenStudy (anonymous):

A boat has a speed of 5mph in calm water. It takes the boat 3 hours to travel upstream it only 2 hours to travel the same distance downstream. What is the value of c. The speed of the current? Will medal for best answer.

OpenStudy (anonymous):

2(5+c)=3(5-c) solve for c 10+2c=15-3c 5c=5 c=1

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

Do you understand how he did that? We know that the distance is the same. So we find d=vt for upstream and downstream when going downstream the current will help the velocity so you add c to the velocity when going upstream the current will work against the velocity of the boat, so you subtract c from the velocity then you just set the equations equal to each other (because the distance is the same for both of them) and then solve for c

OpenStudy (anonymous):

Yep :) I redid it and checked it to make sure.